Impact: 08:50 – 08:55, 09:50 – 09.55
Problem: Upstream/ISP router rebooted. Suspected DDOS. Operator is restarting services and restoring connectivity.
Future resolution: Replacing upstream router.
We’ve now omitted the ISP with connectivity issues and re-routed traffic to different ISP.
Fault: Optical fiber outage causing network disruption of internet traffic.
Status: Partially remedied (failover path in use).
Work is being done to find the fault and remedy it.
Datastorage is not affected by the outage (multipathing).
At 10:40 backup and primary paths are now recovered. Full redundancy has been restored.
Further investigation into the failed optical fiber will be made.
Work is being undertaken to improve and speed up the fail-over of fiber redundancy for inter-DC connectivity.
At 09.00 a switch failure occurred in core infrastructure which burnt optical fiber transceivers.
Remediation: By 09.15 the fault had been found and all network paths had been migrated to secondary fiber connections.
The faulty switch has been replaced and redundancy restored. We’re planning to perform maintenance work (future scheduled window) to implement new protocols for automated fail-over of fiber paths.
Uppströmsleverantör hade problem med sitt nätverk. Det drabbade vissa delar av Adminors nät.
Orsak var DDOS mot leverantör.